OpenStudy (anonymous):

how do i write this equation in slope intercept form? 3x - 2y = 5

OpenStudy (precal):

solve for y

OpenStudy (anonymous):

set the equation to y

OpenStudy (mendicant_bias):

So, slope intercept form is \[y = mx +b\]where m is the slope, and b is the y-intercept. From what you have now, all you have to do is rearrange. A good place to start would be isolating the term with y on it by putting it on one side by itself, right?\[3x - 2y = 5 |||||||||3x - 5 = 2y\]Then, you can divide both sides by the coefficient on y to get it in standard form.\[3x-5=2y|||||||||\frac{ 3}{ 2 }x - \frac{ 5 }{ 2 } = y\]All you need to do now is just flip it to get\[y = \frac{ 3 }{ 2 }x - \frac{ 5 }{ 2 } |||OR||| y = 1.5x - 2.5\]

OpenStudy (mendicant_bias):

*slope-intercept form, NOT standard form, my bad.

OpenStudy (mendicant_bias):

Does this help?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (mendicant_bias):

No problem! Glad I could help. And for future reference, when a question is answered, you should close it. It causes less confusion; some people actually go way back to unanswered problems buried in the list to make sure people get help, but often they're just things that got figured out which weren't closed.

OpenStudy (dwelsher):

y=3/2x-5/2

OpenStudy (dwelsher):

y=mx+b